# XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

### v1

@@ 7.4 Case-3: New Algorithms and Data
Secondly, newly filtered data is continuously incorporated into the LLM for training. XPUTimer has successfully diagnosed anomalies arising from variance in the training data. In one specific case, the algorithm team attempts to train a Llama-80B model with data containing a sequence length of 64k, while the original training script is designed for sequence lengths of 4k. XPUTimer identifies a significant anomalous decline in MFU ($41\%$) on 512 H800 GPUs, accompanied by an increase in $V_{inter}$.
After routing this anomaly to the infrastructure team, the root cause was identified in the dataloader, specifically in the attention mask generation process. When the sequence length is short, the latency incurred by mask generation is minimal. However, the complexity of mask generation scales as $O(L^{2})$, where $L$ represents the sequence length. As a result, the dataloader experiences extremely poor performance when the sequence length increases to 64k.
8 Lessons Learned and Future Work
In this section, we share the practical lessons learned during the development and deployment of XPUTimer, along with directions for future work.
@@ 8.1 Practical Usages
Algorithm teams.
Algorithm teams are dedicated to continuously integrating new innovations into the existing LLM training pipeline. They employ XPUTimer to verify that submitted training jobs meet the expected training throughput. By leveraging XPUTimer, algorithm teams can identify and resolve slowdown anomalies caused by elementary inefficient code, without requiring intervention from the infrastructure team.
Infrastructure team.
The infrastructure team is committed to optimizing GPU kernel libraries, training frameworks, and parallel backbones. With XPUTimer’s lightweight logging system, the infrastructure team can gather sufficient runtime data to analyze submitted jobs and identify new optimization opportunities. Anomalies caused by codebase modifications that cannot be resolved by the algorithm team are ultimately routed to the infrastructure team.
Operations team.
The operations team is responsible for maintaining the stability of all low-level resources. When a training job is terminated due to low-level issues, the operations team queries intra-kernel tracing information provided by XPUTimer. This allows them to identify the faulty machine without requiring a fully comprehensive test.
8.2 Holistic Diagnostics
The training stack for advancing AIGC is inherently complex. Previously, in the absence of a holistic diagnostic framework like XPUTimer, significant effort was wasted on communication among algorithm, infrastructure, and operations teams. Algorithm teams often required collaboration with the infrastructure team to address slowdowns caused by minor but inefficient codebase modifications. When low-level issues arose, the operations team had to conduct extensive low-level benchmarking tests to identify the problem, resulting in delays to the algorithm teams’ training jobs. The holistic diagnostics provided by XPUTimer streamline this process by pinpointing elementary inefficiencies in the codebase for algorithm teams, only routing complex anomalies with aggregated metrics to the infrastructure team, and supplying the operations team with valuable runtime low-level data.
@@ 8.3 Using Historical Data
Historical traces are essential for enhancing the effectiveness of anomaly detection. While runtime data is crucial, it is insufficient on its own for accurately identifying anomalies. XPUTimer detects various issues by comparing real-time data against historical data. For instance, when using issue latency distributions to detect kernel-issue stalls, XPUTimer relies on historical data from specific backbones operating on specific hardware. By collecting historical data from healthy jobs, XPUTimer can efficiently identify anomalies in newly submitted LLM training jobs. In the future, as deployment data grows, we aim to release relevant datasets to further streamline the diagnosis of LLM training.
8.4 Hardware Extensibility
Currently, XPUTimer does support the extensibility of additional hardware, particularly NPUs dedicated to DL training. Its tracing mechanism is not only backbone-agnostic but also designed for seamless extension to other NPUs. Since XPUTimer directly instruments key code segments at the Python and C++ runtime levels, extending it is straightforward, provided the relevant computation kernel is supplied by the NPU vendor.
8.5 Diagnosing Communication Slowdown
Diagnosing communication slowdowns caused by issues such as network jitter is challenging, as runtime tracing can introduce significant overhead or lack accuracy. Currently, XPUTimer utilizes a method similar to MegaScale [2], leveraging NCCL tests to identify faulty machines or switches. We are in the process of developing an eBPF-based [49] tracing tool to accurately monitor bandwidth across RDMA NICs with minimal overhead. In the future, XPUTimer will be further enhanced with these fine-grained tracing capabilities.
@@ 9 Related Work
Anomaly diagnosis in large-scale distributed deep learning training tasks has consistently been a hot research focus[2, 14, 3, 50, 51, 52]. Megascale [2] only focuses on LLM training tasks based on Megatron-LM. It identifies network-related hardware and software issues in the training process by conducting intra-host network tests and NCCL tests. Falcon [14] detects prolonged iterations using the Bayesian Online Change-Point Detection algorithm. C4D [3] modifies the Collective Communication Library to collect message statistics, such as sizes and durations of transfers, to identify the performance bottlenecks. However, these approaches primarily address communication-related performance issues and fail to encompass the anomalies across the LLM training stack. Additionally, they depend heavily on a single parallel backbone, Megatron, which limits their generality. Meanwhile, many researches[50, 51, 52] focus on task recovery. These approaches are orthogonal to XPUTimer, which could seamlessly integrate with XPUTimer.
@@
Many research efforts[53, 54, 55, 56, 57] focus on anomaly diagnosis at different levels of the datacenter, including runtime, network, and storage. SCALENE[53] introduces an algorithm to assist Python programmers in optimizing their code by distinguishing between inefficient Python execution and efficient native execution. AND[54] is a unified application-network diagnosing system that leverages a single metric, TCP retransmissions (TCP retx), to identify network anomalies in cloud-native scenarios. AAsclepius[55] proposes a PathDebugging technique to trace fault linkages between the middle network and autonomous systems. Researchers [56] from Alibaba analyze four key factors that impact SSD failure correlations: drive models, lithography, age, and capacity. As these works address anomaly problems in specific scenarios, they are unable to resolve the challenges targeted by XPUTimer in large-scale distributed LLM training.
@@ 10 Conclusion
In this paper, we introduce XPUTimer, a real-time diagnostic framework for LLM training. By addressing critical challenges such as lightweight long-term monitoring, root cause detection, and backbone extensibility, XPUTimer provides a comprehensive solution for anomaly diagnostics across large-scale GPU clusters. With its novel intra-kernel tracing mechanism and holistic aggregated metrics, XPUTimer not only reduces diagnostic complexity but also improves the detection and resolution of non-obvious slowdowns and errors. Its deployment in real-world training clusters, spanning over 6,000 GPUs, demonstrates its efficacy in diagnosing distributed LLM training. Furthermore, XPUTimer is open-sourced to encourage broader adoption.

### v2

@@ 7.3.3 Case-3: New Algorithms and Data
8 Experience
@@ 8.1 Practical Usages
By leveraging Flare, algorithm teams can independently identify and resolve regression anomalies caused by inefficient code without requiring intervention from the infrastructure team. With Flare’s lightweight logging, the infrastructure team also collects sufficient runtime data to analyze submitted jobs and independently discover new optimization opportunities. Only regression anomalies that cannot be resolved by algorithm teams are routed to the infrastructure team. Overall, the frequency of collaboration on recurrent regressions, such as unnecessary synchronizations, decreased by 63.5% within a one-week deployment, accelerating model evolution for algorithm teams and allowing the infrastructure team to focus more on system optimization.
@@ 8.2 Using Historical Data
Historical traces are essential for enhancing the effectiveness of regression detection. While runtime data is crucial, it is insufficient on its own for accurate regression diagnosis. Flare detects issues by comparing real-time data against historical data. E.g., when using issue latency distributions to detect kernel-issue stalls, Flare relies on historical data from specific backends operating on specific hardware in Figure 11. However, regression detection is inherently difficult to ascertain, even with historical data. Therefore, in practice, we adopt a conservative policy: we report potential regressions and their root causes without directly terminating training jobs.
8.3 Hardware Extending
Currently, Flare has supported the extensibility of additional hardware, particularly NPUs dedicated to DL training. Since Flare directly instruments key code segments at the Python and C++ runtime levels, extending it is straightforward. We already support internal CUDA-native NPUs, incurring less than $0.5\%$ runtime overhead when running on 450 NPUs. We also investigate the extensibility of Flare’s intra-kernel inspecting on NPUs. Our study shows that the methodology is largely extensible, as NPUs also use dedicated hardware cores for cross-device communication. Inspecting their register status provides additional information for the operations team to address hang-error anomalies, but this requires additional engineering effort.
8.4 Scopes & Limitations
Flare operates within certain scopes and has limitations.
Firstly, Flare targets large-scale private GPU clusters for LLM training and relies on historical data to calibrate its diagnostic metrics. This assumption is reasonable. Accessing historical data is practical in a private GPU clusters. In our internal cluster, we also profiled typical LLMs and parallel backends at different scales to obtain ground-truth data for building the diagnostic mechanism.
Secondly, Flare cannot detect regressions in training jobs with major architecture updates, e.g., switching from a dense LLM to a sparse MoE LLM. In such cases, Flare must collect new historical data to refine its diagnostic engine. This requirement is usually reasonable, as Flare is designed to relieve the infrastructure team from repetitive efforts on recurrent regression anomalies, allowing them to focus instead on optimizing LLMs with new architectures.
@@ 9 Related Work
Anomaly diagnosis in large-scale distributed deep learning training tasks has consistently been a hot research focus[jiangMegaScaleScaling, wuGREYHOUNDHunting, dongEnhancingLargescale, peng2018optimus, haider2011fault, chen2020elastic]. Megascale [jiangMegaScaleScaling] only focuses on LLM training tasks based on Megatron-LM. It identifies network-related hardware and software issues in the training process by conducting intra-host network tests and NCCL tests. Greyhound [wuGREYHOUNDHunting] detects prolonged iterations using the Bayesian Online Change-Point Detection algorithm. C4D [dongEnhancingLargescale] modifies the Collective Communication Library to collect message statistics, such as sizes and durations of transfers, to identify the performance bottlenecks. However, these approaches primarily address communication-related performance issues and fail to encompass the anomalies across the LLM training stack. Additionally, they depend heavily on a single parallel backend, limiting their generality.
@@
Many research efforts[bergerTriangulatingPython, wangDiagnosingApplicationnetwork, yangAAsclepiusMonitoring, han2021depth, wangZeroOverhead] focus on anomaly diagnosis at different levels of the datacenter, including runtime, network, and storage. AND[wangDiagnosingApplicationnetwork] is a unified application-network diagnosing system that leverages a single metric, TCP retransmissions (TCP retx), to identify network anomalies in cloud-native scenarios. AAsclepius[yangAAsclepiusMonitoring] proposes a PathDebugging technique to trace fault linkages between the middle network and autonomous systems. Researchers [han2021depth] from Alibaba analyze four key factors that impact SSD failure correlations: drive models, lithography, age, and capacity. As these works address anomaly problems in specific scenarios, they are unable to resolve the challenges targeted by Flare in large-scale distributed LLM training.
@@ 10 Conclusion
In this paper, we introduce Flare, a real-time diagnostic framework for LLM training. By addressing challenges such as lightweight long-term monitoring, root cause detection, and backend extensibility, Flare provides a comprehensive solution for anomaly diagnostics across large-scale GPU clusters. With its novel intra-kernel inspecting mechanism and holistic aggregated metrics, Flare not only reduces diagnostic complexity but also improves the detection and resolution of non-obvious slowdowns and errors. Its deployment in real-world training clusters, spanning over 6,000 GPUs, demonstrates its efficacy in diagnosing distributed LLM training.
Acknowledgments
This work is partially sponsored by the National Key Research and Development Program of China (2024YFB4505700), National Natural Science Foundation of China (62232011) and Natural Science Foundation of Shanghai Municipality (24ZR1430500) and CAAI-Ant Group Research Fund. We thank the anonymous reviewers and our shepherd Andrew Moore for their constructive feedback and suggestions.